Louisiana State University Health Sciences Center Administration & Finance
 

Friday, July 14, 2006

  • Reminder: The boilers in the MEB will be shut down at 5:00 PM on Friday, July 14th so that the building steam distribution valves can be changed out early Saturday morning. The steam will be shut off at 5:00 PM on Friday, July 14th and turned back on at about 6:00 PM on Saturday, July 15th. From Friday, July 14th at 5:00 PM until Saturday, July 15th at 6:00 PM, the MEB domestic hot water will not be hot, and the steam serving the autoclaves in the MEB and in the CSRB on Floors 3-7 will be off.
  • The MDL repairs in the MEB are largely completed. The contractor needs to caulk the joint between the new tile floor and the walls and the tile floor needs to be waxed.
  • The epoxy flooring will be installed over the weekend in the Gross Anatomy Lab, the Swing Lab and the Morgue.
    • An additional vinyl coating may be required in the Morgue to provide additional protection from higher concentrations of certain chemicals.
  • The new casework for the Gross Anatomy Lab, the Swing Lab and the Morgue should be delivered to the site on Monday and installation will begin then. The contractor will also install the emergency shower in the Morgue at the same time that the casework is installed.
  • Progress continues in the two Large Lecture Rooms in the MEB.
    • Carpet is complete in Lecture Room B and almost complete in Lecture Room A.
    • The pedestals to support the seating and table tops have been delivered.
      • The installation of the pedestals is 95+% complete in Lecture Room B.
      • The installation had not started in Lecture Room A, while the carpet installers finished.
    • The installation of the vinyl wallpaper is approximately 50% in B.
    • The network cable has been pulled and bundled beneath the floor in B.
    • The electrical wiring has been pulled in both A and B and it is ready to be run to the table tops when they have been installed.
    • The installation of the new presentation equipment will begin next week.
  • The contractors continue to progress in completing the temporary repairs to the Dental School.
    • The piping of the chiller is approximately 50% complete.
    • The air compressors have been piped in and the electrical connections are completed. The controls need to be programmed and the air dryers need to be tied into the system.
    • The deionized water system has been piped in to the building and the electrical connections are complete.
    • The ice water system and wet vacuum system are scheduled to be installed next week.